Microwave-Microwave Double-Resonance Assignment Of The Gauche Conformer Of Methyl Vinyl Ether

\underline{David A. McWhorter}, Sam Cupp, Chung Yi Lee and Brooks H. Pate, Department of Chemistry, University of Virginia, Charlottesville, VA 22901.

An electric-resonance optothermal spectrometer has been used to perform microwave-microwave doubleresonance measurements on the less stable conformer of methyl vinyl ether. The measured rotational constants are consistent with a gauche structure. The dipole moment components for this conformer have been measured using the Autler-Townes splitting of states that occcurs in the presence of a strong, resonant microwave field. Rotational constants, dipole moments and the threefold barrier to internal rotation, V_{3}, will be reported. We have searched in double-resonance for the second tunneling component for the equivalent gauche conformer but have found no transitions. This suggests the tunneling gap is larger than about 2GHz.